OpenStudy (anonymous):

Ln (x+1) - Ln (x-1) = 3

OpenStudy (anonymous):

Use log properties and combine ln[(x+1)/(x-1)] = 3 raise both sides by e (x+1) / (x-1) = e^3 x+1 = e^3(x-1) x+1 = xe^3 - e^3 x-xe^3 = -e^3 - 1 x(1-e^3) = -e^3-1 x = (-e^3-1) / (1-e^3) Have a good day!

OpenStudy (anonymous):

note that lna-lnb = ln(a/b) and lne^x=x ln(x+1)-ln(x-1)=3 ln((x+1)/(x-1))=lne^3 ((x+1)/(x-1))= e^3 ((x+1)/(x-1))(x-1) = e^3 (x-1) x+1 = (x-1)e^3 = xe^3 -e^3 x+1-xe^3=xe^3-e^3 -xe^3 x-xe^3+1-1=-e^3 -1 (1-e^3)x=-e^3-1 x=(-e^3-1)/(1-e^3)

OpenStudy (anonymous):

Sorry that I am a bit slow

jhonyy9 (jhonyy9):

so what is equal continued x= -(e^3 +1)/(-(e^3 -1) = (e^3 +1)/(e^3 -1)

jhonyy9 (jhonyy9):

hope you like it !!!

jhonyy9 (jhonyy9):

yes i know but this is the final result

OpenStudy (lgbasallote):

oh wait...thought expanding to cubey thingy makes difference..seems not